OpenStudy (anonymous):

A polynomial has been factored, as shown below: f(x) = (x − 6)(x + 5)(x − 9) What are the zeros of the polynomial? 6, −5, 9 6, 5, 9 −6, 5, −9 −6, 6, −5, 5, −9, 9

Vocaloid (vocaloid):

split the function into three parts and solve (x-6) = 0 (x+5) = 0 (x-9) = 0 solve for x

OpenStudy (anonymous):

A: 6, −5, 9 B: 6, 5, 9 C: −6, 5, −9 D: −6, 6, −5, 5, −9, 9

OpenStudy (anonymous):

i dont really get it

Vocaloid (vocaloid):

(x-6) = 0 what is x?

OpenStudy (anonymous):

what do you mean what is x?

Vocaloid (vocaloid):

what is the value of x?

Vocaloid (vocaloid):

x = a number, what is that number?

Vocaloid (vocaloid):

x - 6 = 0 x = ?

OpenStudy (anonymous):

-6

Vocaloid (vocaloid):

nope, try one more time x - 6 = 0 x = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

i dont really get it sorry

Vocaloid (vocaloid):

yup! now we move to the next part: (x+5) = 0 x = ?

OpenStudy (anonymous):

oh yay

OpenStudy (anonymous):

-5?

Vocaloid (vocaloid):

yup! next part: (x − 9) = 0 x = ?

OpenStudy (anonymous):

9

Vocaloid (vocaloid):

yup, so our zeros are x = 6, x = 9, and x = -5

OpenStudy (anonymous):

thank you i got it i gave you a medal and ima fan right now